OpenStudy (anonymous):

Write your answer as a coordinate pair. y=2x-4 x+y=5

OpenStudy (anonymous):

x+ (2x-4) = 5 3x - 4 =5 3x = 9 x=3

OpenStudy (anonymous):

3+y = 5 y= 2 (3,2)